Shift Summary: 6/2 Owl Shift

Miguel Barrio (mbarrio@hep.uchicago.edu)
Wed, 2 Jun 1999 07:34:00 -0500

Shift Summary for Owl shift, 06/02/99

On Shift: Julie Whitmore
Miguel Barrio

Others present:
Sasha Glazov working on SODs and shtuffs pertaining the DCs.
Val's voice on the speaker yelling "Tiiiime to panic, the
eerrrrrr-m-s is too wide, caaaall the CsI expert..."
Ed and Rick came in around 4:00 when we started to have beam.

Hours of beam delivered at all intensities: 3.0
Hours of beam delivered at nominal intensity: 0.0
Hours of 99 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): ?

Accounting of lost time:

Got beam since 4:00. System hung up a couple of times. No useful data
collected.

Activities:

0:30 Ramped magnet to Ptkick = 205 MeV/c (799 configuration).
0:33 Start run 13285 (BEAM_832_ECAL config).
1:00 Control room calls and says we'd have beam in ~30 minutes.
4:00 Started to get beam (not very good quality. apparently).
Beam profile looks bad, long tails.
MCR notified that they don't have access to swic info from
switchyard (us, not US).
4:12 Run 13285 dies, gateway problem (ddd vetos, plns 1,2 enabled)
DC plots don't look good at all in spite of having modified
the prescales (B01 set to 20, B07 set to 20).
6:20 Gas checked ok.
6:30 Viv called due to multiple gateway hangups. Dart ppl were
notified.
7:10 Called the control room. They are filling just ONE batch
which means that our instanteneous intensity is about
4E12.

=========================================================================
Runs ended this shift:

13285: died, no data flushed.
13286 thru 13290: all died due to gateway problems. Some of these runs'
online histos were saved.

Tapes brought to FCC this shift: 0

-- 
Miguel Barrio
Office: HEP 319 (773-702-7484)
Email:  mbarrio@hep.uchicago.edu